OpenStudy (thats_kyy):

An unknown number y is 15 more than an unknown number x. The number y is also x less than 8. The equations to find x and y are shown below. y = x + 15 y = −x + 8 Which of the following statements is a correct step to find x and y?

OpenStudy (anonymous):

y = x + 15 y = 8 - x i think they're both right

OpenStudy (mathstudent55):

Of course both equations are correct. They were given with the problem, and we're told they are the equations needed to solve the problem.

OpenStudy (thats_kyy):

Multiply The Equations to eliminate y Add The Equations To Eliminate x Write the points where the graph of the equations intersect the x-axis

OpenStudy (mathstudent55):

The second equation comes from y is x less than 8 That is y = 8 - x which can be rewritten as y = -x + 8

OpenStudy (mathstudent55):

The equations are correct. This is a given.

OpenStudy (thats_kyy):

But whats The Answer?

OpenStudy (mathstudent55):

The question is choosing the next step in solving this problem from a list of choices. We can only help if we see the choices.

OpenStudy (thats_kyy):

okay ill put it up

OpenStudy (mathstudent55):

Ok.

OpenStudy (thats_kyy):

Multiply the equations to eliminate y Add the equations to eliminate Write the points where the graphs of equations intersect the x-axis write the points where the graphs of equation intersect the y - axis @mathstudent55

OpenStudy (mathstudent55):

Notice that in the first equation you have x, and in the second equation you have -x. If you add x to -x, you get zero. If you add the equations, x will be eliminated. Then all you have is the variable y which you can solve for. It seems like adding the equations is a good step in solving this system of equations.

OpenStudy (thats_kyy):

Okay

OpenStudy (mathstudent55):

Answer is B.

OpenStudy (thats_kyy):

Thank You!!

OpenStudy (mathstudent55):

You're welcome.